OpenStudy (anonymous):

find the second derivatives of y= x^3(2x+4)^2 a. 2x+4 b. x^3(2x+80) c. x^2(80x+192) d. x^2(20x+60)

OpenStudy (anonymous):

i find it x^2(4x+8)

OpenStudy (anonymous):

But the answer you chose it's not listed on the multiple choices.

OpenStudy (anonymous):

yes thats y i think the answer is e for error LOL

OpenStudy (anonymous):

do you find it the same with me?

OpenStudy (anonymous):

I feel like it's A or B.

OpenStudy (anonymous):

letter a has a chance w8 i check it

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

product rule... twice.. what a pain (go on wolf)

OpenStudy (bahrom7893):

woahh wait...

OpenStudy (bahrom7893):

meh.. ill just solve it by hand

OpenStudy (anonymous):

theres a 20x in your wolf i see the light from letter d XD

OpenStudy (bahrom7893):

still simplifyin the 2nd derivative....

OpenStudy (bahrom7893):

i cant factor it into any of the forms..., neither can wolframalpha :/ are u sure the question's correct?

OpenStudy (anonymous):

so the real answer is finally shown letter E =error thank u bah

OpenStudy (bahrom7893):

lol np.. well at least my answer was right.. haha 16x(5x^2+12x+6), wolf confirmed it.

OpenStudy (anonymous):

so you get the 2nd derivati easy enough..can i have a tip on derivative..do you think i need to memorize all the formula so i can get it easy on quiz? i have quiz on derivative tommorow

OpenStudy (anonymous):

\[(fg)'=f'g+g'f\] so in this case you would use \[f(x)=x^2,f'(x)=2x\] \[g(x)=(2x+4)^2,g'(x)=2(2x+4)\times 2=4(2x+4)=8x+16\]

OpenStudy (anonymous):

oww the f o g so your answer to is letter e? XD